OpenStudy (anonymous):

"The hypotenuse of right triangle is 2 units longer than its height. Determine a formula for the height as a function of the base b.

OpenStudy (mathteacher1729):

Have you tried drawing a picture of the triangle and its lengths?

OpenStudy (anonymous):

So far I have h^2 + b^2 = (h+2)^2

OpenStudy (anonymous):

\[h = \sqrt{c^{2} - b^{2}},\], where c = h + 2 is the hypotenuse. Then \[h = \sqrt{(h + 2)^{2} - b^{2}}\] \[h^{2} = h^{2}+ 4h + 4 - b^{2}\]\[4h = b^{2} - 4\] \[h = (b^{2} - 4)/4.\]

OpenStudy (anonymous):

Oh! I see. I was solving for the wrong variable. Thanks!

OpenStudy (anonymous):

:^)